# LLM-Enhanced Static Analysis for Precise Identification of Vulnerable OSS Versions

### v1

@@ 8 MXFContext *mxf = arg;
15-  if (item_num > 65536) {
16+  if (item_num > 65536 || item_num < 0) {
@@
Example 1) A vulnerability in FFmpeg (CVE-2017-14169 (CVE, 2017b)) shows the limitation of existing approaches in terms of capturing vulnerability logic. The code snippet of pc is shown in Listing LABEL:list:CVE-2017-14169. A sanitizing check was added at Line 16 for the item_num. Without this check, a remote attacker can make a crafted file with a large item_num field such as 0xffffffff, causing a buffer overflow issue in the avio_read function at Line 28 and potentially leads the application to exhibit incorrect behavior or crash.
V-SZZ (Bao et al., 2022) assumes the deletion lines in the security patch as vulnerable codes and the basic idea of it is to pinpoint the earliest commit introducing the deletion lines as vic. However, vulnerability logic is composed of various vulnerable codes, such as Lines 24 and 28 in Listing LABEL:list:CVE-2017-14169. V-SZZ only considers tracing back deleted line (Line 15) and overlooks the real vulnerability behavior, leading to incorrect identification of vic. V0Finder (Woo et al., 2021) attempts to identify vulnerable versions using code clone detection. It generates hash values for the entire vulnerability function and identifies code clones according to the distance value of two hashes. This method may still introduce excessive vulnerability-unrelated features due to the inclusion of “entire function”.
@@
2index 57fbcd7a1b341..bab62a08c1c1d 100644
3--- a/libavformat/avidec.c
4$+$ b/libavformat/avidec.c
5@@ -350,8 +350,7 @@ static void avi_read_nikon(AVFormatContext *s, uint64_t end)
@@
10- if (avio_tell(s->pb) + size > tag_end)
11-    size = tag_end - avio_tell(s->pb);
12+ size = FFMIN(size, tag_end - avio_tell
13+    (s->pb));
@@
Listing 2: Motivating Example of Semantic-Invariant Modification in Refactoring Commit.
@@
Example 2) After extracting the vulnerability logic and its corresponding vulnerable statements ($S_{v}$), we can locate the vic by backtracing the code changes of $S_{v}$ in the previous commit, thereby identifying the vulnerable versions. V-SZZ (Bao et al., 2022) identifies code clones before and after such code changes by the edit distance. However, refactoring commits, prevalent in OSS development, aims to improve and reorganize code while preserving its functionality. Such commits can result in situations where the code’s semantics remain the same, but there is a significant syntactic difference between the code before and after modifications. Essentially, such a commit would cause the tool like V-SZZ to incorrectly flag it out as a vic.
To exemplify such a scenario, we present code changes made in the file avidec.c of FFmpeg. As depicted in Listing LABEL:semantic, the added lines convey identical semantic meaning with deleted lines, which both represent selecting the minimum value between size and tag_end-avio_tell(s->pb) and assigning it to the variable size.
Due to the refactoring of the if structure into the FFMIN function, the edit distance utilized by V-SZZ (Bao et al., 2022) perceives a significant syntactic difference between the two (48% similarity), leading to incorrect pinpoint of the vic. ReDeBug (Jang et al., 2012) and V0Finder (Woo et al., 2021) relies on syntax-based pattern matching for code clone detection. MVP (Xiao et al., 2020) and AFV (Shi et al., 2022) extract the dependencies between two statements and normalize them to hash values as function signatures, overlooking the semantic information of statements.
Observation. This case emphasizes the significance of identifying semantic invariance in the code modifications, which was often overlooked in existing techniques. AST is a tree representation of code that explicit statement order and the execution logic, which contains rich semantic information (Yang et al., 2021). Vercation introduces a code clone detection grounded in expanded AST analysis.
Figure 2. Vercation Workflow.
3. Design of Vercation
@@
We propose Vercation, an end-to-end automated approach designed for vulnerable version identification in OSS vulnerabilities. The high-level workflow is illustrated in Figure 2. Vercation consists of three phases: Vulnerable code extraction (P1), Code changes detection (P2) and Vulnerable version range delineation (P3).
@@
In P1, Vercation combines program slicing and LLM to identify and extract vulnerability-related program statements in a precise manner. Specifically, we utilize the patch code as a slicing criterion to extract dangerous flows. To mitigate the risk of including statements unrelated to the vulnerability (false positives), we employ prompt construction with few-shot and Chain-of-Thought (CoT) strategies, empowering LLM to reason with the vulnerability based on the extracted dangerous flows and accurately extract vulnerability-related statements.
@@
In P2, Vercation retraces historical commits to collect previous modifications of vulnerable statements. For each statement before and after modification, Vercation expands the functions within the statement, generates ASTs and normalizes them. Then we utilize an in-order traversal algorithm to compare the AST before and after the modification as semantic similarity, which determines if the commit is the initial introduction of the vulnerable statements.
In P3, Vercation identifies the versions affected by a vulnerability based on the CVE’s pc and the corresponding vic.
In the following sections, we explain these phases in detail.
3.1. Vulnerable Code Extraction
@@
Taking a vulnerability patch commit as an input, Vercation identifies and extracts program statements related to vulnerability in three steps as explained in the next three subsections respectively.
3.1.1. Dangerous flow extraction
@@
Firstly, Vercation conducts program slicing (Weiser, 1984) on the source code to extract crucial program statements that may contribute to the vulnerability, which we call them as dangerous flow. The slicing is based on the following criteria: deleted/added statements and patch-related variables. Building upon existing techniques (Xiao et al., 2020; Shi et al., 2022; Woo et al., 2022), we execute program slicing on the program dependency graph of the patched function according to the slicing criteria. For different statement types, we customize different slicing rules:
@@
  • Assignment statement affects the data-flow values of vulnerability behavior. We conduct normal slicing in the assignment statements and the assigned variables should be added to the slicing criterion. For example, if we take variable item_num of Line 15 and 16 in Listing LABEL:list:CVE-2017-14169 as slicing criterion and perform forward slicing, Line 24 and 27 are included.
@@
  • Conditional statement affects the reaching condition of vulnerability trigger. We aim to slice the condition statement that takes the variables of slicing criterion as condition check, and the result also includes all the subsequent statements of condition statement (e.g., subsequent statement Line 17-18 of condition statement Line 16, subsequent statement Line 26 of condition statement Line 25).
@@
  • Return statement. There is no need for forward slicing because there is no dependency between the return value and the statements following the return statement. For example, there is no need for forward slicing on Line 18 and 26 in Listing LABEL:list:CVE-2017-14169.
Meanwhile, the slicing is performed in two directions: backward and forward. Backward slicing is to trace the source of patch-related variables. For example, when we designate item_num (e.g., Line 15, 16 in Listing LABEL:list:CVE-2017-14169) as the criterion, the result of the backward slice is Line 9 (Line 15 is data-dependent on Line 9). Additionally, Line 15 is control-dependent on Line 11, as its execution is contingent on the unsatisfied condition in Line 11; otherwise, the function will promptly return. Following backward slicing, we retain Lines 9 and 11 as backward dangerous flow. Forward slicing is used to find the trigger behavior that causes the vulnerability. For example, in Listing LABEL:list:CVE-2017-14169, if we set the modified statements Lines 15 and 16 along with the variable item_num as criteria, the outcomes of forward slicing include 1) Lines 24, 27 and 28, which are directly or indirectly data-dependent on the variable item_num and 2) Lines 17, 18, 25 and 26, which are directly or indirectly control-dependent on the variable item_num.
3.1.2. Dangerous flow-based prompt construction
@@
Secondly, we leverage an LLM to refine these dangerous flows further since the dangerous flows extracted in the above step (program slicing) may include false positives (unrelated to vulnerability logic). We explore prompting strategies to assist LLM in vulnerability comprehension task and in refining vulnerable statements from the dangerous flows extracted through program slicing.
@@
Our method utilizes an LLM that supports interaction with system prompts and user prompts. The system prompt sets the role and background of the LLM. The user prompt consists of specific instructions issued by the user. At the beginning of the system prompt, we assign the role of a security researcher to the model with the statement, ”You are a security researcher, an expert in detecting security vulnerabilities,” and indicate that we will provide the CVE information and dangerous flow. The system prompt concludes with a declaration of the fixed structure expected for the model’s response. The content of the user prompt includes detailed CVE information including CVE ID, CWE ID, CVE description, and dangerous flow extracted from program slicing.
Additionally, we use the following strategies to determine the features of the user prompt:
1) Few-shot prompting. Few-shot prompting is a technique where we provide the LLM with a small number of examples demonstrating the desired task before presenting the actual problem. This approach helps guide the model’s behavior by showing it how to respond to similar inputs. We include two carefully crafted examples in the user prompt before providing the target CVE information. Each example consists of: a) A sample CVE ID, CWE ID and CVE description. b) Corresponding dangerous flow extracted from program slicing. c) The expected vulnerability logic analysis. d) The correctly identified vulnerable statements.
2) Chain-of-thought prompting. Chain-of-thought (CoT) prompting provides LLM with a prompt that encourages it to generate intermediate reasoning steps before arriving at a final answer (Wei et al., 2022). We prompt the model to ”Please reason the vulnerability logic from the provided code” before concluding the vulnerable statements.
Our prompting template is shown in Figure 3 and we take the vulnerable statements in the answer as the input of the next step.
System:You are a security researcher, expert in detecting security vulnerabilities.I will provide you with a CWE ID, CWE ID, CVE description, and dangerous code.Please extract the vulnerability logic from the code and indicate which statementsare relevant to the vulnerability logic.Provide a response only in the following format: vulnerability logic: ‹text¿vulnerable lines : [¡Line number List¿]Do not include the added line number (with +) and anything else in response.User:¡CVE ID¿,¡CWE ID¿,¡CVE description¿ and ‹dangerous flows¿
Figure 3. The LLM Prompt Template.
3.1.3. Statement weight allocation
As vulnerable statements contribute differently to vulnerability triggers, we propose strategies to allocate weights to the vulnerable statements extracted by LLM. We have noticed that certain types of C/C++ vulnerability often involve specific functions. For instance, in Listing 1, Lines 24 and 28 are critical trigger lines. With this knowledge, we can rely on a pre-defined table (Table 1) to identify such functions within vulnerable statements extracted by LLM. We summarized the common types of vulnerabilities in C/C++, and combined the authors’ domain knowledge with the enumeration of sensitive functions from relevant papers (Shu et al., 2017; Alaba et al., 2017) to create the pre-defined table.
Some OSS commonly used customized functions serve as variants of sensitive functions. For instance, the av_calloc function on Line 24 is a variant of the calloc function. To discern whether these customized functions contain sensitive functions, we conduct inter-procedure analysis. Taking Listing LABEL:list:CVE-2017-14169 as an example, considering this is an overflow-type vulnerability, we recognize the av_calloc in Line 24 and avio_read function in line 28 as sensitive functions.
Parameters. There are two parameters for the weight allocation of statements, i.e., weight_v and weight_d, that could affect the effectiveness of Vercation: 1) weight_d is the default weight for initializing the weights of all statements. 2) weight_v is the weight for the statements containing dangerous functions. We set weight_v=2$\times$×weight_d, weight_d=1.0. Note that these weights are used in Code changes detection phase (P2).
3.2. Code Clone Detection
Given the vulnerable statements $S_{v}$ of a patch commit (extracted in Section 3.1), Vercation performs a commit backtrace. The goal is to identify which previous commit introduced $S_{v}$ by tracing the history of earlier modifications. To achieve this, Vercation employs the git blame command to backtrack through previous commits. We introduce a semantic-level similarity comparison method to automatically detect code clones before and after modifications. We define the statements of the post-modification commit as $S$, and those from the pre-modification commit as $S^{\prime}$, with $S_{v}\subseteq S$.
@@ 3.2. Code Clone Detection
We first use line mapping to pre-filter highly similar lines in $S^{\prime}$ with $S_{v}$. If such lines are found, it indicates that the commit is not the first to introduce $S_{v}$. We use the edit distance to calculate line similarity $Similarity_{A}(S_{v},S_{i}^{{}^{\prime}})$, $S_{i}\in S^{\prime}$. We set a similarity threshold $\vartheta_{1}$, if $Similarity_{A}(S_{v},S_{i}^{{}^{\prime}})\geq\vartheta_{1}$, we continue to backtrack the previous commit. If $Similarity_{A}(S_{v},S_{i}^{{}^{\prime}})<\vartheta_{1}$, we proceed to make a comparison with the next semantic matching.
Table 1. Sensitive Functions for Some Vulnerability Types.

### v2

@@ 5 MXFContext *mxf = arg;
12-  if (item_num > 65536) {
13+  if (item_num > 65536 $|$ item_num < 0) {
@@
2@@ -350,8 +350,7 @@ static void avi_read_nikon(AVFormatContext *s, uint64_t end)
@@
7- if (avio_tell(s->pb) + size > tag_end)
8-    size = tag_end - avio_tell(s->pb);
9+ size = FFMIN(size, tag_end - avio_tell
10+    (s->pb));
@@
13
14/*The definition of method FFMIN in libavutil/common.h*/
15#define FFMIN(a, b) ((a)>(b)?(b):(a))
Listing 2: Motivating Example of Code Refactoring Commit.
@@
Example 2) After extracting the vulnerability logic and its corresponding vulnerable statements ($S_{v}$), we can locate the vic by backtracing the code changes of $S_{v}$ in the previous commit, thereby identifying the vulnerable versions. V-SZZ [17] identifies code clones before and after such code changes by the edit distance. However, refactoring commits, prevalent in OSS development, aim to optimize and reorganize code while preserving its functionality [42, 43].
A refactoring commit in FFmpeg (Listing LABEL:semantic) demonstrates how structural changes caused by method encapsulation mislead existing detection methods. In this commit, developers extracted the code with if condition structure into a new method FFMIN() to improve code modularity. While this refactoring preserves functionality, it introduces significant syntactic divergence between the original code (Lines 7-8) and the restructured code (Lines 9-10).
Traditional tools like V-SZZ [17] consider edit distance as the metric to detect code clones. The edit distance between the original and refactored code drops to 48% due to the method outlining, causing V-SZZ to incorrectly flag this commit as the vulnerability-introducing commit (vic). Other approaches like ReDeBug [14] and V0Finder [16] also fail to recognize the equivalence between inline logic and encapsulated methods, as their coarse pattern matching ignores structural abstraction.
Observation. This case highlights the structural refactoring in OSS development and the limitations of existing clone detection techniques. Vercation addresses this by expanding function calls during AST generation—inlining the FFMIN() method body to reconstruct the original logic. Combined with AST normalization, this allows Vercation to detect code logic equivalence despite structural variations, accurately tracing the vic through refactored commits.
III Design of Vercation
@@
We propose Vercation, an end-to-end automated approach designed for vulnerable version identification in OSS vulnerabilities. The high-level workflow is illustrated in Figure 2. Vercation consists of three phases: Vulnerable code extraction (P1), Code clone detection (P2) and Vulnerable version range determination (P3).
Figure 2: Vercation Workflow.
@@
In P1, Vercation combines program slicing and LLM to identify and extract vulnerability-related program statements in a precise manner. Specifically, we utilize the patch code as a slicing criterion to extract dangerous flows, defined as program statements that directly or indirectly affect the variables or expressions at the patch code. To mitigate the risk of including statements unrelated to the vulnerability (false positives), we employ prompt engineering with few-shot and Chain-of-Thought (CoT) strategies, empowering LLM to reason with the vulnerability based on the extracted dangerous flows and accurately extract vulnerability-related statements.
@@
In P2, Vercation retraces historical commits to collect previous modifications of vulnerable statements. For each statement before and after modification, Vercation expands the functions within the statement, generates ASTs, and normalizes them. Then we utilize an in-order traversal algorithm to compare the AST before and after the modification as code similarity, which determines if the commit is the initial introduction of the vulnerable statements.
In P3, Vercation identifies the affected versions based on the CVE’s vfc and the corresponding vic.
To note, Vercation supports patches that span multiple functions and files. In P1, we extract dangerous flows from each affected function independently, then combine all extracted dangerous flows as input to the LLM for unified vulnerability logic analysis. In P2, we perform a backtrack analysis for vulnerable statements from all affected functions.
III-A Vulnerable Code Extraction
@@
Taking a vfc as an input, Vercation identifies and extracts program statements related to a vulnerability in three steps, as explained in the following subsections respectively.
III-A1 Dangerous Flow Extraction
@@
Firstly, Vercation conducts program slicing [44] on the source code to extract crucial program statements that may contribute to the vulnerability, which we call dangerous flow. The slicing is based on the following criteria: deleted/added statements and patch-related variables. Building upon existing techniques [12, 13, 15], we execute program slicing on the program dependency graph of the patched function according to the slicing criteria. We performed slicing in two directions: backward and forward. Backward slicing is used to trace the source of patch-related variables and forward slicing is used to find the trigger behavior that causes the vulnerability. For different statement types, we customize different slicing rules:
@@
  • Assignment statement affects the data-flow values of vulnerability behavior. We conduct normal slicing in the assignment statements and the assigned variables should be added to the slicing criterion. For example, if we take item_num of Line 12 and 13 in Listing LABEL:list:CVE-2017-14169 as slicing criterion and perform forward slicing, Line 21 and 24 are included.
@@
  • Conditional statement affects the reaching condition of the vulnerability trigger. We aim to slice the condition statement that takes the variables of slicing criterion as condition check, and the result also includes all the subsequent statements of condition statement (e.g., subsequent statement Line 14-15 of condition statement Line 13, subsequent statement Line 23 of condition statement Line 22).
@@
  • Return statement. There is no need for forward slicing because there is no dependency between the return value and the statements following the return statement. For example, there is no need for forward slicing on Line 15 and 23 in Listing LABEL:list:CVE-2017-14169.
III-A2 Dangerous Flow-based Prompt Construction
@@
Secondly, we leverage an LLM to refine these dangerous flows further since the dangerous flows extracted in the above step (program slicing) may include false positives (unrelated to vulnerability logic). We explore prompting strategies to assist LLM in the vulnerability comprehension task and in refining vulnerable statements from the dangerous flows extracted through program slicing. It is important to note that the LLM’s role is to perform code understanding and vulnerability logic analysis rather than to recall specific CVE information from its training data. Most vulnerability reports do not provide detailed vulnerability logic analysis or identify specific vulnerable code statements [45], where the LLM’s semantic analysis capabilities become essential.
@@
Our method utilizes an LLM that supports interaction with system prompts and user prompts. The system prompt sets the role and background of the LLM. The user prompt consists of specific instructions issued by the user. At the beginning of the system prompt, we assign the role of a security researcher to the model with the statement, “You are a security researcher an expert in detecting security vulnerabilities,” and indicate that we will provide the CVE information and dangerous flow. The prompt also concludes with a declaration of the fixed output format expected for the model’s response. The content of the user prompt includes detailed CVE information including CVE ID, CWE ID, CVE description, and dangerous flow extracted from program slicing. The extracted dangerous flows are formatted with line numbers prefixed to each statement when constructing prompts, such as “4442 total_size += msec->size; 4444 stash->info_ptr_memory = (bfd_byte *) bfd_malloc (total_size);”.
Additionally, we use the following prompt strategies to optimize the performance of LLM:
  1. (i)
    Few-shot prompting is a technique where we provide the LLM with a small number of examples demonstrating the desired task before presenting the actual problem. We carefully select two examples from publicly available and verified CVE cases. To avoid data leakage, the cases are not included in our experimental dataset. The selection of examples follows specific criteria: (1) having complete and clear CVE descriptions, (2) containing explicit vulnerability fix code, and (3) featuring easily understandable vulnerability logic. These examples were created and reviewed through a systematic process: analyzing CVE descriptions and patch code, extracting key vulnerability logic, annotating relevant vulnerable statements, and writing clear vulnerability analysis explanations. The examples were created by the first author and reviewed by the fourth and fifth authors. To ensure experimental reproducibility, we maintain a fixed set of examples throughout all experiments. Each example consists of: a) A sample CVE ID, CWE ID and CVE description, b) Corresponding dangerous flow extracted from program slicing, c) The expected vulnerability logic analysis, and d) The correctly identified vulnerable statements.
  2. (ii)
    Chain-of-thought (CoT) prompting provides LLM with a prompt that encourages it to generate intermediate reasoning steps before arriving at a final answer [46]. We prompt the model to “Please analyze the code following these steps: 1. Explain the vulnerability logic from the code 2. Indicate which statements are relevant to the vulnerability logic”. This structured format naturally implements a chain of thought - requiring the model to first reason about and explain the vulnerability logic (the reasoning step) before identifying the specific vulnerable lines (the identification step). This approach identifies the vulnerable statements using the code understanding capability of LLM rather than superficial pattern matching. Table 3 shows our prompting template, and we take the vulnerable statements from the response as the input for the next step.
Figure 3: The LLM Prompt Structure.
III-B Code Clone Detection
Given the vulnerable statements $S_{v}$ of a fixing commit (extracted in Section III-A), Vercation performs a commit backtrack to identify which previous commit introduced $S_{v}$ by tracing the history of earlier modifications. To achieve this, Vercation employs the git blame command to backtrack through previous commits. Figure 4 illustrates our code clone detection process. For each intermediate commit between the vfc and the potential vic, we compare the post-modification statements ($S$) with the pre-modification statements ($S^{\prime}$) using two levels of similarity detection: a quick syntactic check using edit distance and a deeper structural analysis using ASTs. The results from both comparisons, combined with statement weights assigned in Section III-B3, contribute to a final similarity score. If this score is below the threshold $\theta_{3}$, we identify the commit as the vic. Otherwise, we continue the backtrack process.
III-B1 Syntactic Similarity Analysis
@@
First, we use line mapping to pre-filter highly similar lines in $S^{\prime}$ with $S$, where $S\subseteq S_{v}$. If such lines are found, it indicates that the commit is not the first to introduce $S_{v}$. We use the edit distance to calculate line similarity $\mathrm{Similarity}_{A}(S^{\prime}_{i},S_{i})$, where $S^{\prime}_{i}\in S^{\prime}$ and $S_{i}\in S$. We set a similarity threshold $\vartheta_{1}$; if $\mathrm{Similarity}_{A}(S^{\prime}_{i},S_{i})\geq\vartheta_{1}$, we continue to backtrack the previous commit. If $\mathrm{Similarity}_{A}(S^{\prime}_{i},S_{i})<\vartheta_{1}$, we proceed to the more detailed structural analysis. To note, when encountering merge commits with multiple parents during backtrack, we traverse each parent path independently and select the chronologically earliest vic across all branches.
Figure 4: Overview of Code Clone Detection Process in Vercation.
III-B2 Structural Similarity Analysis
As exemplified in Example 2 of Section II-C, Lines 5-6 (deleted pre-commit lines) and 7-8 (added line of post-commit, e.g., included in $S_{v}$) exhibit high similarity in code behavior. However, previous research failed to identify this high similarity through simple line mapping, resulting in the commit (version) being erroneously labeled as an introducing commit (vulnerable version). We propose a fine-grained similarity comparison method at the AST-level to address this limitation. AST is a tree representation of code that preserves well-defined components of statements, explicit statement order and the execution logic [47].
Method Expansion. Due to code refactoring often outlining code into a new method, as shown in Listing LABEL:semantic, the codes in Lines 7-8 are outlined into methods FFMIN. The standard AST comparison treats the if-condition (Lines 7-8) as completely different from the FFMIN function call, resulting in a false vulnerability identification. Therefore, we utilize the inline technique to expand the method during AST generation to gather more code behavior within the statements. This involves obtaining the method definition at the callsite and expanding the method body inline, resulting in a more comprehensive code representation. We get the AST for the target file using Clang without performing actual compilation. Furthermore, through recursive retrieval of line numbers for sub-ASTs, we can locate the sub-AST corresponding to $S_{v}$.
AST Normalization. To address structural variations arising from different developer coding styles, it is essential to normalize the AST. This process mitigates inconsistencies and reduces analytical noise by standardizing code representations. AST normalization involves conditional structure normalization, loop structure normalization, and relational operation normalization.
  • Loop structure can be represented as while, do-while and for structures, we normalize all the loop structure to while [48]. The for structure is for(initialize; condition; increment) {statements;}, we transform it into while structure: initialize; while(condition) {statements; increment}.
  • Relational operation include commutative and noncommutative operation. Commutative relational operation (i.e., +, !=, &&) refers to the operation where commutate the operands does not change the operation semantics. We normalize the two operands (left and right child nodes) of the operation (parent nodes) by alphabetical order. For example, we transform $b+a$ (b is the left child node) to $a+b$ (a is the left child node). Noncommutative relational operation (i.e., >>, <=<=) will change the operation semantic when commutating the operands, we normalize them by predefined rules. For example, we transform $b<a$ to $a>b$ (normalize operation << to >>).
AST Similarity. Ultimately, we utilize an in-order traversal to ascertain whether the commit introduces $S$ by comparing the AST similarity of $S_{i}$ and $S^{\prime}_{i}$. First, we perform an in-order traversal of the ASTs for $S_{i}$ and $S^{\prime}_{i}$, converting the results into sequences $Sq_{i}$ and $S^{\prime}q_{i}$. Then, we calculate the edit distance between the two sequences as their AST-based similarity $\mathrm{Similarity}_{B}(Sq_{i},S^{\prime}q_{i})$. We set a similarity threshold $\vartheta_{2}$, and if $\mathrm{Similarity}_{B}(Sq_{i},S^{\prime}q_{i})\geq\vartheta_{2}$, $S_{i}$ and $S^{\prime}_{i}$ are considered to be AST-based similar. If $\mathrm{Similarity}_{B}(Sq_{i},S^{\prime}q_{i})<\vartheta_{2}$, it indicates that there is no statement similar to $S_{i}$ in the pre-modification commit. During this process, we classify a commit as a refactoring commit when it exhibits low edit-distance similarity ($\mathrm{Similarity}_{A}(S_{i},S^{\prime}_{i})<\vartheta_{1}$) but high AST-based similarity ($\mathrm{Similarity}_{B}(Sq_{i},S^{\prime}q_{i})\geq\vartheta_{2}$). This indicates that while the code structure has been significantly modified, the underlying functionality remains unchanged.
III-B3 Statement Weight Allocation
Not all statements identified by an LLM contribute equally to vulnerability in triggering the exploit. To address this, we propose a weight allocation strategy that prioritizes vulnerable statements based on their invocation of known sensitive functions. The core premise is that the presence of specific, high-risk functions within a statement is a strong indicator of its criticality. We followed a systematic process to identify and categorize these sensitive functions. Starting with a comprehensive review of previous studies [49, 50], we collected existing sensitive function enumerations. We then conducted a preliminary study of common C/C++ vulnerabilities to identify the most frequent vulnerability categories, such as buffer overflow, integer overflow, and use-after-free, and examined the functions commonly involved in known vulnerabilities for each type. This process involved three authors with security expertise: the first author proposed the initial categorization, which was then independently reviewed and validated by the fourth and fifth authors, with any discrepancies resolved through discussion. The result is the pre-defined table (Table I) of sensitive functions for various vulnerability types. We analyze each vulnerable statement extracted by the LLM. If a statement contains a direct call to any function listed in our sensitive function table, it is assigned a higher weight, marking it as a probable vulnerability trigger. For instance, in Listing LABEL:list:CVE-2017-14169, Lines 21 and 25 are critical trigger lines.
Furthermore, we conduct an inter-procedural analysis to identify customized functions that serve as variants of canonical sensitive functions. For instance, the av_calloc function on Line 21 is a variant of the calloc function. To discern whether the customized function F contains sensitive functions, we first analyze the definition and implementation code of function F, then check if F’s function body calls any sensitive functions from Table I. If such call relationships are found, F is also identified as a sensitive function. Taking Listing LABEL:list:CVE-2017-14169 as an example, considering this is an overflow-type vulnerability, we recognize the av_calloc in Line 21 and avio_read function in line 25 as sensitive functions.
TABLE I: Sensitive Functions for Some Vulnerability Types.